OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = -5x + 2, solve for f(g(x)) when x = 3

jigglypuff314 (jigglypuff314):

so first plug in 3 into the function of g(x)

OpenStudy (anonymous):

so would I get f(3)=7 and g(3)=-13 ?

jigglypuff314 (jigglypuff314):

yes for the g(x) part but f(g(x)) means to take what you got from g(x) and plug that into the function of f(x)

OpenStudy (anonymous):

oh ok so after i got the g(3)=-13 where would i plug it into the f(x)?

jigglypuff314 (jigglypuff314):

f(x) = 2x + 1 f(g(x)) = 2(-13) + 1

OpenStudy (anonymous):

If i had the question The function for the cost of materials to make a hat is f(x) = x + 1, where x is the number of hats. The function for the selling price of those hats is g(f(x)), where g(x) = x + 2. Find the selling price of 10 hats.. would i do the same thing?

jigglypuff314 (jigglypuff314):

if you are finding g(f(x)) then you would plug in 10 into the f(x) function first but basically the same idea so yes :)

OpenStudy (anonymous):

thank you :)